A principal of $500 is invested in an account at 7% per year compounded annually. What is the total amount of money in the account after 5 years?
Answer by stanbon(75887)

A principal of $500 is invested in an account at 7% per year compounded annually. What is the total amount of money in the account after 5 years?
----------
A(t) = P(1+(r/n))^(nt)
-------
A(5) = 500(1+(0.07/1))^(1*5)
-------
A(5) = 500(1.07)^5
A(5)= $701.28
====================
Cheers,
Stan H.
====================